The Basics of Capital Budgeting: Payback

Payback period was the earliest -Select-capital structurefinancial statementcapital budgetingCorrect 1 of Item 1 selection criterion. The -Select-NPVMIRRIRRpaybackCorrect 2 of Item 1 is a "break-even" calculation in the sense that if a project's cash flows come in at the expected rate, the project will break even. The equation is:

The -Select-shorterlongerCorrect 3 of Item 1 a project's payback, the better the project is. However, payback has 3 main disadvantages: (1) All dollars received in different years are given -Select-lessequalmoreCorrect 4 of Item 1 weight. (2) Cash flows beyond the payback year are ignored. (3) The payback merely indicates when a project's investment will be recovered. There is no necessary relationship between a given payback and investor wealth maximization.

A variant of the regular payback is the discounted payback. Unlike regular payback, the discounted payback considers -Select-projectcapitaloverheadCorrect 5 of Item 1 costs. However, the discounted payback still disregards cash flows -Select-duringbeforebeyondCorrect 6 of Item 1 the payback year. In addition, there is no specific payback rule to justify project acceptance. Both methods provide information about -Select-profitabilitywealthliquidityCorrect 7 of Item 1 and risk.

Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 7%.

0 1 2 3 4
Project A -1,050 650 395 220 270
Project B -1,050 250 330 370 720

What is Project A's payback? Do not round intermediate calculations. Round your answer to four decimal places.

years

What is Project A's discounted payback? Do not round intermediate calculations. Round your answer to four decimal places.

years

What is Project B's payback? Do not round intermediate calculations. Round your answer to four decimal places.

years

What is Project B's discounted payback? Do not round intermediate calculations. Round your answer to four decimal places.

years
Answer #1

DS fx -SUM(D3:07) C B D 1 Project A year pv@7% Cash flows (1,050.00) $ 650.00 $ 395.00 220.00 $ 270.00 Present value Cumulati

f =SUM(D3:07) с DS А 1 Project A 2 year 30 pv@7% Cash flows -1050 650 395 =C3/1.07 =C4/1.07 =C5/1.07 =C6/1.07 NPV Present val